OpenStudy (anonymous):

Solve the equation for x. If necessary, round your answer to two decimal places. Do not include "x =" in your answer. e^x=2

OpenStudy (anonymous):

is x=2

OpenStudy (decarr432):

Is this the entire problem?

OpenStudy (jhannybean):

\[e^x = 2\]\[\ln e^x = \ln(2)\]\[x = \ln(2)\]